OpenStudy (mikeymaximum):

I shall give a medal 2 whoever gives me the answer 2 each of my Qs and if i get a good score then I'll fan every1 that had helped :)

OpenStudy (mikeymaximum):

Starting with this 1 :)

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

winx club!!

OpenStudy (anonymous):

yes

OpenStudy (mikeymaximum):

Solve the equation. x +1/6= 6 A.x =–6 1/6 B.x =–5 5/6 C.x =5 5/6 D.x =6 1/6

OpenStudy (mikeymaximum):

Don't mind the ?s

OpenStudy (anonymous):

C im sure

OpenStudy (anonymous):

5 5/6

OpenStudy (anonymous):

please medal me

OpenStudy (anonymous):

sure

OpenStudy (mikeymaximum):

NEW Q :D

OpenStudy (anonymous):

yay can u fan me?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thx daphne

OpenStudy (bohotness):

:)

OpenStudy (anonymous):

hello @MikeyMaximum

OpenStudy (mikeymaximum):

I posted it

OpenStudy (mikeymaximum):

I posted it as a different Q